OpenStudy (anonymous):

Can someone show me step by step how to do this? Using points (2,5) and (5,7)... A) Find the standard form of the equation of the line. B) Find the slope-intercept form of the equation of the line.

OpenStudy (anonymous):

first try to plot and connect two point with line then determine the slope of that line next define point at x=0 for y intercept and using the same slope to determine y

OpenStudy (anonymous):



OpenStudy (anonymous):

first find the value of m i.e m = y2-y1/x2-x1 then the equation of line is y-y2=m(x-x2) then put the values of m, y2 and x2 and u will get the equation.... try it

OpenStudy (anonymous):

create another triangle which connect to the y intercep, composed of x=0 and the same slope with the first one you'll get